Circular No. 6637

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 F. Frontera, Università di Ferrara and Istituto Tecnologie e Studio Ra-

diazioni Extraterrestri (ITESRE), Consiglio Nazionale Ricerche, Bologna; J. Greiner, Astrophysikalisches Institut, Potsdam; L. A. Antonelli, BeppoSAX, Rome; D. Dal Fiume and M. Orlandini, ITESRE; T. Boller and W. Voges, Max-Planck-Institut für Extraterrestrische Physik, Garching; E. Costa, M. Feroci and L. Piro, Istituto Astrofisica Spaziale, Frascati; and G. Zavattini, Università di Ferrara, communicate: “A 34 280-s ROSAT HRI observation centered at the BeppoSAX/WFC position of the γ-ray burst GRB 970228 (IAUC 6572), performed between Mar. 10.7875 and 13.32 UT, detected eight sources in the HRI field-of-view, of which only one is in the 3′ WFC error radius and within the 50′′ error box of the BeppoSAX fading source SAX J0501.7+1146 (IAUC 6576). The position of the pointlike source, RX J050146+1146.8, has the following coordinates: α = 5^{h}01^{m}46^{s}.7, δ = +11^{o}46′51′′ (equinox 2000.0), with an error radius of 10′′, coinciding (within 2′′) with the optical transient (IAUC 6584). Using only the 10351 s of low background level, the source countrate is (1.0 ± 0.3) × 10^{−3} count/s. Assuming a power-law spectrum with photon index 1.9, the corresponding flux in the band 0.1–2.4 keV is (3.8 ± 1.2) × 10^{−14} erg cm^{−2} s^{−1}, uncorrected for any galactic absorption.”

GRO J1655–40 E. Kuulkers, University of Oxford; T. Belloni, M. Mendez, M. van der

Klis, and R. Wijnands, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam (UoA); and J. van Paradijs, University of Alabama at Huntsville and UoA, report: “Using RXTE All Sky Monitor data, we discovered about twenty drops in the 2–12-keV intensity by ∼ 25–95 percent during 90-s measurements of the relativistic jet blackhole binary GRO J1655–40 in its current outburst (IAUC 6393). During these drops, the spectral hardness increases. Public RXTE PCA observations on Feb. 26.8– 27.0 UT show two additional drops by ∼ 92 percent with durations of order 1 min. A period of 2.621 ± 0.001 days (1σ) is best-fit with the occurrence of the ASM and PCA intensity drops. This is consistent with the optically determined period of the system (Orosz and Bailyn 1997, Ap.J. 477, 876). All drops occurred between orbital phases 0.72 and 0.86. This constitutes the first orbital evidence of GRO J1655–40 in x-rays. Similar orbital dips have previously been seen in low-mass x-ray binary ‘dip’ sources. Further monitoring at all wavelengths at these orbital phases is encouraged.”

1997 April 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green